# Feature Space Analysis by Guided Diffusion Model

### v1

@@ Appendix A $10$ Carefully-chosen Tusker Images
Appendix B Effect of $w_{g}$
To check the effect of $w_{g}$ in Eq. (3), this experiment aims to investigate a transition of squared Euclidean distances between features of generated images and the ones of their corresponding actual images on $w_{g}$ values of $0.5$, 11, 22, 44 and $8$. Specifically, using CLIP’s image encoder with ResNet-50 backbone as the target feature extractor, $\boldsymbol{f}_{s}$ is defined as the feature of each of the $15$ actual images in Figs. 3 and 4. Except $w_{g}$, all the other hyper-parameters are fixed as configured in Section III-B. Similar to the experiments before, this experiment targets the best image whose feature yields the shortest squared Euclidean distance to $\boldsymbol{f}_{s}$ in the image generation process.
Fig. 10 depicts a box plot that summarises the median, $25$ and $75$ percentiles, and the maximum and minimum of squared Euclidean distances between the features of the $15$ best images generated by each $w_{g}$ value and the ones of the corresponding actual images. As seen from Fig. 10, $w_{g}=8$ leads to images whose features exhibit shorter squared Euclidean distances than $w_{g}=4$ used in Section IV. Regarding this, Figs. 11 and 12 presents the images generated by each $w_{g}$ value. As can be seen from the top row and the second row from the bottom of Fig. 11 and the second row from the bottom of Fig. 12, $w_{g}=8$ sometimes results in too unrealistic images despite the short squared Euclidean distances. This can hinder the feature space analysis by visually checking generated images. Thus, $w_{g}=4$ is adopted to balance between minimising squared Euclidean distances and maintaining the realness of generated images.
@@ Appendix B Effect of $w_{g}$
Appendix C Effect of the Early Step Emphasis Strategy
Using $w_{g}=4$ in Appendix B, an ablation study is conducted to compare image generation approaches with and without the early step emphasis strategy. That is, the former uses $K=1000$ and $K=8$ self-recurrence iterations at the first $T^{\prime}=5$ reverse steps and the others, respectively, whereas the latter uses $K=8$ at all steps. When using the early step emphasis strategy, the average squared Euclidean distance between the features of the $15$ generated images in Figs. 11 and 12 and the ones of the corresponding actual images is $0.06554\pm 0.03468$. In contrast, without this strategy, the average is $0.06906\pm 0.03185$. This verifies the effectiveness of the early step emphasis strategy for generating images whose features match user-specified ones more closely.
Appendix D Effect of $\nabla_{thres}$
To demonstrate $\nabla_{thres}$’s effect on our decoder’s image generation, Fig. 13 shows images generated by setting $\nabla_{thres}$ to three times the standard deviation of values in $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$ when using ViT-H/14 as the target feature extractor. As seen from Fig. 13, a generated image often contains an artifactual (unintended) large black region that is caused by Eq. (3) where the predicted source noise $\epsilon_{\theta}\left(\boldsymbol{z}_{t},t\right)$ is excessively modified with large gradient values in $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$. The maximum gradient value used in this modification is reduced using a smaller $\nabla_{thres}$ like twice the standard deviation of values in $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$, as is the case for Fig. 7. As a result, a generated image becomes less likely to contain an artifactual large black region, although several images in Fig. 7 still have such regions. In addition, setting $\nabla_{thres}$ to a further lower value prevents our decoder from generating images whose features are sufficiently close to a user-specified feature. One possible solution for artifactual large black regions is discussed in Section IV-D.

### v2

@@ 6 Declaration of Generative AI and AI-assisted Technologies in the Manuscript Preparation Process
Part B Effect of $w_{g}$
To check the effect of $w_{g}$ in Eq. (3), this experiment aims to investigate a transition of squared Euclidean distances between features of generated images and the ones of their corresponding actual images on $w_{g}$ values of $0.5$, 11, 22, 44 and $8$. Specifically, using CLIP’s image encoder with ResNet-50 backbone as the target feature extractor, $\boldsymbol{f}_{s}$ is defined as the feature of each of the $15$ actual images in Figs. 3 and 4. Except $w_{g}$, all the other hyper-parameters are fixed as configured in Section 3.2. Similar to the experiments before, this experiment targets the best image whose feature yields the shortest squared Euclidean distance to $\boldsymbol{f}_{s}$ in the image generation process.
Fig. 10 depicts a box plot that summarises the median, $25$ and $75$ percentiles, and the maximum and minimum of squared Euclidean distances between the features of the $15$ best images generated by each $w_{g}$ value and the ones of the corresponding actual images. As seen from Fig. 10, $w_{g}=8$ leads to images whose features exhibit shorter squared Euclidean distances than $w_{g}=4$ used in Section 4. Regarding this, Figs. 11 and 12 presents the images generated by each $w_{g}$ value. As can be seen from the top row and the second row from the bottom of Fig. 11 and the second row from the bottom of Fig. 12, $w_{g}=8$ sometimes results in too unrealistic images despite the short squared Euclidean distances. This can hinder the feature space analysis by visually checking generated images. Thus, $w_{g}=4$ is adopted to balance between minimising squared Euclidean distances and maintaining the realness of generated images.
@@
Part C Effect of the Early Step Emphasis Strategy
Using $w_{g}=4$ in B, an ablation study is conducted to compare image generation approaches with and without the early step emphasis strategy. That is, the former uses $K=1000$ and $K=8$ self-recurrence iterations at the first $T^{\prime}=5$ reverse steps and the others, respectively, whereas the latter uses $K=8$ at all steps. When using the early step emphasis strategy, the average squared Euclidean distance between the features of the $15$ generated images in Figs. 11 and 12 and the ones of the corresponding actual images is $0.06554\pm 0.03468$. In contrast, without this strategy, the average is $0.06906\pm 0.03185$. This verifies the effectiveness of the early step emphasis strategy for generating images whose features match user-specified ones more closely.
Part D Effect of $\nabla_{thres}$
To demonstrate $\nabla_{thres}$’s effect on our decoder’s image generation, Fig. 13 shows images generated by setting $\nabla_{thres}$ to three times the standard deviation of values in $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$ when using ViT-H/14 as the target feature extractor. As seen from Fig. 13, a generated image often contains an artifactual (unintended) large black region that is caused by Eq. (3) where the predicted source noise $\epsilon_{\theta}\left(\boldsymbol{z}_{t},t\right)$ is excessively modified with large gradient values in $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$. The maximum gradient value used in this modification is reduced using a smaller $\nabla_{thres}$ like twice the standard deviation of values in $\nabla_{\boldsymbol{z}_{t}}l\left(f\left(\hat{\boldsymbol{x}}_{t}\right),\boldsymbol{f}_{s}\right)$, as is the case for Fig. 7. As a result, a generated image becomes less likely to contain an artifactual large black region, although several images in Fig. 7 still have such regions. In addition, setting $\nabla_{thres}$ to a further lower value prevents our decoder from generating images whose features are sufficiently close to a user-specified feature. One possible solution for artifactual large black regions is discussed in Section 4.4.1.